OpenStudy (anonymous):

Write a direct variation equation that relates x and y. Assume that y varies directly as x. Then solve. If y = –5 when x = 15, find y when x = –1.

OpenStudy (anonymous):

a. y=-1/5x;1/5 b. y=-1/3x;2/5 c. y=1/3x;-1/3 d. y=-1/3x;1/3

OpenStudy (e.mccormick):

Which part are you getting stuck on?

OpenStudy (anonymous):

how to do it

OpenStudy (e.mccormick):

Direct variation means there is some sort of ratio between the values, and it is always the same ratio. So, they start with \(y=-5\) and \(x=15\). Do you see a ratio between those numbers?